I've done this problem so many dfferent way that all I have done is confuse myself more and more. Please someone help me!
The question is: Determine the equation of the line. Put in slope-intercept form, if possible.
Through (2,8), perpendicular to 5x+8y=74
A) y=8/5x
B) y=5/8x-24
C) -8/5x-24/5
D) y=8/5x+24/5
I'm thinking the answer is A, but I'm not real sure about it.

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)

Put the given equation into slope-intercept form by solving for y. The coefficient on the x term will then be the slope. Lines are perpendicular if and only if their slopes are negative reciprocals. Only one of the given answers has a slope that is the negative reciprocal of the slope of the given line.

Answer by stanbon(75887)

Determine the equation of the line. Put in slope-intercept form, if possible.
Through (2,8), perpendicular to 5x+8y=74
----------
The given line has slope = -5/8
So the line you want must have slope 8/5 and must pass thru (2,8)
The form of the equation is y = mx + b
You know y=8 when x=2 and you know m=(8/5); solve for "b":
8 = 2(8/5) + b
b = (40/5)-16/5 = 24/5
--------------------
EQUATION:
y = (8/5)x + (24/5)
---------------

=======================
Cheers,
Stan H.